# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ 5.2. The Construction of Degree-Aware Star Substructure Synopses (DAS3)
Based on the degree grouping, we will construct $m$ degree-aware star substructure synopses (DAS3), $Syn_{j}$ (for $1\leq j\leq m$), which are essentially $m$ grid files (Güting, 1994), respectively, that store (aggregated) vertex embeddings for star substructures in specific degree groups.
@@
  • a list of MBRs, $v_{i}.MBR_{\delta}$ (for $\delta\in(\delta_{j-1},\delta_{j}]$), that minimally bound all embedding vectors, $o^{\prime}(s_{v_{i}})$, of star substructures $s_{v_{i}}$ with degrees equal to $\delta$.
To facilitate the access of cells in DAS3 synopsis $Syn_{j}$ efficiently, we sort all cells in descending order of their keys $C.key$. Here, for each cell $C$, we use $C.UB$ to denote its maximal corner point (i.e., taking the maximum value of the cell interval on each dimension). The key, $C.key$, is defined as $||C.UB||_{2}$, where $||Z||_{2}=\sum_{i=1}^{d}Z[i]^{2}$ for a $d$-dimensional vector $Z$.
Note that, if a query vertex embedding $o(q_{i})$ dominates $C.UB$, then the key $||o(q_{i})||_{2}$ of query vertex embedding $o(q_{i})$ must be smaller than or equal to the cell key $C.key$. Thus, with such a sorted list of cells, we can efficiently access all cells that are fully or partially dominated by a given query vertex embedding $o(q_{i})$.
@@
Moreover, we also associate corner point $v_{i}.UB_{ub\_\delta}$ with the list of MBRs, $v_{i}.MBR_{\delta}$ (for each $\delta\in(\delta_{j-1},ub\_\delta]$), in synopsis $Syn_{j}$, which can be used for further pruning for a specific degree in interval $(\delta_{j-1},\delta_{j}]$.
Discussions on MBR Computations and Incremental Maintenance: Note that, the MBRs, $v_{i}.MBR_{\delta}$, mentioned above minimally bound embedding vectors $o^{\prime}(v_{i})$ for all star substructures $s_{v_{i}}$. Since there are an exponential number of possible star structures (i.e., $2^{deg(v_{i})}$), it is not efficient to enumerate all star substructures, compute their embedding vectors, and obtain the MBRs. Therefore, in this paper, we propose a sorting-based method to obtain $v_{i}.MBR_{\delta}$ without enumerating all star substructures with degree $\delta$.
Specifically, for each vertex $v_{i}$, we maintain $d$ sorted lists, each of which, $v_{i}.spur\_list_{k}$ (for $1\leq k\leq d$), contains the $k$-th elements, $y_{j}[k]$, of SPUR vectors $y_{j}$ (obtained from $v_{i}$’s 1-hop neighbors $v_{j}\in\mathcal{N}(v_{i})$) in ascending order.
@@
(5) $v_{i}.MBR_{\delta}[2k]$ $=$ $\sum_{deg\in[1,\delta]}v_{i}.spur\_list_{k}[deg],$
(6) $v_{i}.MBR_{\delta}[2k+1]$ $=$ $\sum_{deg\in[deg(v_{i})-\delta+1,deg(v_{i})]}v_{i}.spur\_list_{k}[deg],$
@@
Intuitively, Eqs. (5) and (6) give the lower/upper bounds of the MBR $v_{i}.MBR_{\delta}$ on the $k$-th dimension, by summing up the first $\delta$ and the last $\delta$ values, respectively, in the sorted list $v_{i}.spur\_list_{k}$.
@@
6. DSM Query Processing
In this section, we will illustrate query processing algorithms based on vertex dominance embeddings in DAS3 synopses to efficiently answer DSM queries.
@@ 6.1. Synopsis Pruning Strategies
Moreover, if $o^{\prime}(q_{i})$ does not dominate embedding upper bound vector $v_{i}.UB_{\delta}$, i.e., $v_{i}.UB_{\delta}\notin DR(o^{\prime}(q_{i}))$, then $o^{\prime}(q_{i})$ does not dominate any star substructure $s_{v_{i}}$ with center vertex $v_{i}$ and the corresponding degree group. In other words, $q_{i}$ does not match with $v_{i}$. Thus, vertex $v_{i}$ can be safely pruned. ∎
@@
2compute vertex dominance embeddings $o^{\prime}(q_{i})$ for all query vertices $q_{i}\in V(q)$
@@ 19 $A(q,t)\leftarrow\emptyset$
20 invoke Refinement$(q,Q,G_{t},A(q,t),\emptyset,0)$ to obtain actual DSM query answers in $A(q,t)$
@@
Algorithm 2 Initial DSM Answer Set Generation
6.2. Initial DSM Answer Set Generation
@@
The Synopsis Search: For each query vertex $q_{i}\in V(q)$, we first need to find a synopsis $Syn_{j}$ that matches the degree group of $q_{i}$ (i.e., $deg(q_{i})\in(\delta_{j-1},\delta_{j}]$ must hold; line 3).
As mentioned in Section 5.2, cells, $C$, in the DAS3 synopsis $Syn_{j}$ are sorted in descending order of their keys $C.key$. We thus only need to search for those cells $C\in Syn_{j}$ satisfying the condition that $C.key\geq||o^{\prime}(q_{i})||_{2}$ (line 4). Then, for each candidate vertex $v_{i}\in C.list$, we apply our proposed embedding dominance and MBR range pruning strategies (as discussed in Lemmas 6.1 and 6.2, respectively), and obtain the candidate vertex set, $q_{i}.cand\_set$, for the query vertex $q_{i}$ (lines 5-8).
@@
Input: i) a query graph $q$; ii) a sorted list (query plan) $Q$; iii) a snapshot graph $G_{t}$; iv) matching results set $A$; v) a sorted list, $M$, of vertices matching with query vertices in $Q$, and; vi) the recursion depth $n$
@@ 21       Refinement ($q,Q,G_{t},A,M,n+1$);
Specifically, we maintain a vertex vector $M$ that will store vertices of the subgraph $g$ matching with ordered query vertices in $Q$. To enumerate all matching subgraphs, each time we recursively expand partial matching results by including a new candidate vertex $u$ as $M[n]$ that maps with the $n$-th query vertex $Q[n]$. If we find all vertices in $M$ mapping with $Q$ (i.e., recursion depth $n=|Q|$), then we can add $M$ to the answer set $A$ (lines 1-3). Otherwise, we will find a vertex candidate set $S_{cand}$ from $Q[n].cand\_set$, such that each vertex $u$ in $S_{cand}$ has the same edge connections to that in $M$ as that in $Q$ (lines 5-11). Next, for each vertex $u\in S_{cand}$, we treat it as vertex $M[n]$ matching with $Q[n]$, and recursively call function Refinement($\cdot$) with more depth $n+1$ (lines 13-17).
@@
[…]
@@
6.3. DSM Query Answering
Given a dynamic graph $G_{t-1}$ at timestamp $(t-1)$, a registered query graph $q$, and a graph update operation $\Delta G_{t}$ at timestamp $t$, Algorithm 4 provides a DSM query answering algorithm, which maintains/obtains the latest subgraph matching answer set, $A(q,t)$.
Specifically, if the graph update operation $\Delta G_{t}$ is an insertion of edge $e=(v_{i},v_{j})$, we will calculate an incremental update, $\Delta A(q,t)$, of new subgraph matching results (lines 1-11).
Edge Insertion: For the insertion of an edge $e=(v_{i},v_{j})$, we will first find matching edge(s) $(q_{i},q_{j})$ in the query graph with the same labels (line 3). Then, we will check whether or not edges $(v_{i},v_{j})$ and $(q_{i},q_{j})$ match each other, by applying the embedding dominance and MBR range pruning strategies (described in Section 6.1; lines 4-5). If the answer is yes, then we will generate a new query plan $Q^{\prime}$ starting from query vertices $q_{i}$ and $q_{j}$ (note: the remaining ones are obtained from $Q$, similar to line 10 of Algorithm 2; line 7), initialize the first two matching vertices $v_{i}$ and $v_{j}$ in a sorted list $M^{\prime}$ (lines 8-9), and invoke the function Refinement($\cdot$) with parameters of new query plan $Q^{\prime}$, incremental answer set $\Delta A(q,t)$, the initialized sorted list $M^{\prime}$, and recursion depth $n=2$ (as two matching vertices have been found; line 10). After that, we can update the answer set $A(q,t)$ with $A(q,t-1)\cup\Delta A(q,t)$ (line 11).
Edge Deletion: When $\Delta G_{t}$ is an edge deletion operation $(e-,t)$, we can simply remove those existing subgraph matching answers $g\in A(q,t-1)$ which contain the deleted edge $e$, and obtain the updated answer set $A(q,t)$ (lines 12-16).
Finally, we return the latest DSM query answer set $A(q,t)$ at timestamp $t$ (line 17).
For the edge deletion operation (lines 12-16), since we delete matching results from $A(q,t-1)$ that contain the deleted edge, by using the hash file to check the edge existence with $O(1)$ cost, the time complexity is given by $O(|A(q,t-1)|)$.
Input: i) a dynamic graph $G_{t-1}$ at previous timestamp $(t-1)$; ii) a graph update operation $\Delta G_{t}$; iii) a registered query graph $q$, and; iv) matching result set $A(q,t-1)$ at timestamp $(t-1)$
Output: an updated set, $A(q,t)$, of subgraph matching results at timestamp $t$
@@ 3       $\Delta A(q,t)\leftarrow\emptyset$
8                         $M^{\prime}[0]=v_{i}$
9                         $M^{\prime}[1]=v_{j}$
10                         invoke Refinement$(q,Q^{\prime},G_{t},\Delta A(q,t),M^{\prime},2)$ to incrementally obtain new matching results in $\Delta A(q,t)$
@@ 14      $A(q,t)\leftarrow A(q,t-1)\cup\Delta A(q,t)$
      // edge deletion operation $\Delta G_{t}=(e-,t)$ for edge $e=(v_{i},v_{j})$
17       $A(q,t)=A(q,t-1)$
@@ 20                   $A(q,t)\leftarrow A(q,t)-\{g\}$
Algorithm 4 The Query Answering Algorithm for Continuous Dynamic Subgraph Matching Query

### v3

@@ 5.2. The Construction of Degree-Aware Star Substructure Synopses (DAS3)
Based on the degree grouping $\mathcal{B}=\{B_{1},B_{2},\cdots,B_{m}\}$, we will construct $m$ degree-aware star substructure synopses (DAS3), $Syn_{j}$ (for $1\leq j\leq m$), which are essentially $m$ grid files (Güting, 1994), respectively. In each grid synopsis $Syn_{j}$, we store data vertices in different cells according to their (aggregated) vertex embeddings of star substructures in the corresponding degree group $B_{j}$.
@@
  • a list of MBRs, $v_{i}.MBR_{\delta}$ (for $\delta\in(\delta_{j-1},\delta_{j}]$), that minimally bound all embedding vectors $o^{\prime}(v_{i})$ of star substructures $s_{v_{i}}$ with degrees equal to $\delta$.
To facilitate the access of cells in DAS3 synopsis $Syn_{j}$ efficiently, we sort all cells in descending order of their keys $C.key$. Here, for each cell $C$, we use $C.UB$ to denote its maximal corner point (i.e., taking the maximum value of the cell interval on each dimension). The key, $C.key$, is defined as $||C.UB||_{2}$, where $Z||_{2}=\sum_{i=1}^{d}Z[i]^{2}$ for a $d$-dimensional vector $Z$. Note that, if a query vertex embedding $o(q_{i})$ dominates $C.UB$, then the key $o(q_{i})||_{2}$ of query vertex embedding $o(q_{i})$ must be smaller than or equal to the cell key $C.key$. Thus, with such a sorted list of cells, we can efficiently access all cells that are fully or partially dominated by a given query vertex embedding $o(q_{i})$.
@@
Moreover, we also associate the corner point $v_{i}.UB_{ub\_\delta}$ with the list of MBRs, $v_{i}.MBR_{\delta}$ (for each $\delta\in(\delta_{j-1},ub\_\delta]$), in synopsis $Syn_{j}$, which can be used for further pruning for a specific degree in interval $(\delta_{j-1},\delta_{j}]$.
Figure 9. An example of DAS3 data structure ($\alpha=0.01$, $\beta=100$, and $z_{1}=(0.36,0.64)$).
Discussions on MBR Computations and Incremental Maintenance: Note that, the MBRs, $v_{i}.MBR_{\delta}$, mentioned above minimally bound embedding vectors $o^{\prime}(v_{i})$ for all star substructures $s_{v_{i}}$. Since all possible star structures of $v_{i}$ share the same SPUR vector $x_{i}$ in $o^{\prime}(v_{i})$, we can directly assign the lower and upper bounds of the $k$-th dimension in the MBR (i.e., $v_{i}.MBR_{\delta}[2k]$ and $v_{i}.MBR_{\delta}[2k+1]$, for $0\leq k<d$) as follows:
(5) $\hskip-4.30554ptv_{i}.MBR_{\delta}[2k]=v_{i}.MBR_{\delta}[2k+1]=o^{\prime}(v_{i})[k]=\alpha x_{i}[k]+\beta z_{i}[k].$
For the SPAN vector part in the MBR, since there are an exponential number of possible star structures (i.e., $2^{deg(v_{i})}$), it is not efficient to enumerate all star substructures, compute their embedding vectors, and obtain the MBRs. Therefore, in this paper, we propose a sorting-based method to obtain $v_{i}.MBR_{\delta}$ without enumerating all star substructures with degree $\delta$.
Specifically, for each vertex $v_{i}$, we maintain $d$ sorted lists, each of which, $v_{i}.spur\_list_{k}$ (for $0\leq k<d$), contains the $k$-th elements, $x_{j}[k]$, of SPUR vectors $x_{j}$ (obtained from $v_{i}$’s 1-hop neighbors $v_{j}\in\mathcal{N}(v_{i})$) in ascending order.
@@
(6) $v_{i}.MBR_{\delta}[d+2k]$ $=$ $\sum_{deg\in[1,\delta]}v_{i}.spur\_list_{k}[deg],$
(7) $v_{i}.MBR_{\delta}[d+2k+1]$ $=$ $\hskip-21.52771pt\sum_{deg\in[deg(v_{i})-\delta+1,deg(v_{i})]}\hskip-34.44434ptv_{i}.spur\_list_{k}[deg],$
@@
Intuitively, Eqs. (6) and (7) give the lower/upper bounds of the MBR $v_{i}.MBR_{\delta}$ on the $(d+k)$-th dimension, by summing up the first $\delta$ (smallest) and the last $\delta$ (largest) values, respectively, in the sorted list $v_{i}.spur\_list_{k}$.
@@
The Space Complexity of DAS3 Synopses: Given a data graph $G_{D}$ and the dimension, $d$, of SPUR/SPAN vector, the space complexity of vertex embeddings in Section 4 is $O(|V(G_{D})|\cdot 2d)$. In DAS3, each vertex $v_{i}$ is associated with an embedding upper bound vector $v_{i}.UB_{\delta}$ and a list of MBRs, $v_{i}.MBR_{\delta}$. Moreover, to support incremental updates to vertex embeddings, each vertex maintains $d$ sorted lists $v_{i}.spur\_list_{k}$ (for $0\leq k<d$). Thus, the space complexity of DAS3 synopses is given by $O(|V(G_{D})|\cdot 2d)+\sum_{v_{i}\in V(G_{D})}(deg(v_{i})\cdot 4d+deg(v_{i})\cdot d)$. Since $\sum_{v_{i}\in V(G_{D})}deg(v_{i})=2|E(G_{D})|$, the overall space complexity of our DIVINE approach is given by $O(d\cdot(|V(G_{D})|+|E(G_{D})|))$, which is linear to the graph size.
Example 0.
(Continued with Example 5.1) In the example of Figure 8, we consider two degree groups $B_{1}=(0,1]$ and $B_{2}=(1,3]$. As shown in Figure 9, we construct 2 DAS3 synopses (grids of cells with equal size), $Syn_{1}$ and $Syn_{2}$, for embeddings $o^{\prime}(v_{1})$ of star substructures $s_{v_{1}}^{i}$ falling into degree buckets $B_{1}$ and $B_{2}$, respectively. Specifically, in $Syn_{2}$, we store data vertices $v_{1}$ and $v_{2}$ in the cell $C$, which are associated with embedding upper bound vectors, $v_{1}.UB_{3}$ and $v_{2}.UB_{2}$, and MBR lists, $\{v_{1}.MBR_{2}$, $v_{1}.MBR_{3}\}$ and $\{v_{2}.MBR_{2}\}$, respectively. Here, each MBR $v_{i}.MBR_{\delta}$ (e.g., $v_{1}.MBR_{2}$ in Figure 8) minimally bounds embedding vectors $o^{\prime}(v_{i})$ of star substructures with degree $\delta$ (i.e., $s_{v_{1}}^{4}$, $s_{v_{1}}^{5}$, and $s_{v_{1}}^{6}$ with degree 2), and each embedding upper bound vector $v_{i}.UB_{ub\_\delta}$ (e.g., $v_{1}.UB_{3}$) is the maximum corner point of MBRs in the MBR list (i.e., maximum corner point of MBRs, $v_{1}.MBR_{2}$ and $v_{1}.MBR_{3}$).
6. CSM Query Processing
In this section, we will illustrate query processing algorithms based on vertex dominance embeddings in DAS3 synopses to efficiently answer CSM queries.
@@ 6.1. Synopsis Pruning Strategies
Moreover, if $o^{\prime}(q_{i})$ does not dominate embedding upper bound vector $v_{i}.UB_{\delta}$, i.e., $v_{i}.UB_{\delta}\notin DR(o^{\prime}(q_{i}))$, then $o^{\prime}(q_{i})$ does not dominate any star substructure $s_{v_{i}}$ with center vertex $v_{i}$ and the corresponding degree group. In other words, $q_{i}$ does not match $v_{i}$. Thus, vertex $v_{i}$ can be safely pruned. ∎
@@
2compute vertex dominance embeddings $o^{\prime}(q_{i})$ for all $q_{i}\in V(q)$
@@ 19 $A(q,t)\leftarrow\emptyset$
[…]
@@
Algorithm 2 Initial CSM Answer Set Generation
6.2. Initial CSM Answer Set Generation
@@
[…]
Example 0.
(Continued with Example 5.2) In DAS3 of Figure 9, the query vertex $q_{1}$ with degree 2 will only be compared with data vertices in synopsis $Syn_{2}$ (degree bucket $B_{2}=(1,3]$). As mentioned above, by checking the dominance condition between embedding upper bound vector $q_{1}.UB_{2}$ and the cells’ keys $C.key$, we can obtain candidate cells. Then, for each vertex $v_{i}$ in these cells, we apply our proposed embedding dominance and MBR range pruning strategies (as discussed in Lemmas 6.1 and 6.2, respectively). Specifically, to implement embedding dominance and MBR range pruning strategies, we compare embedding upper bound vectors $q_{1}.UB_{2}$ and $v_{i}.UB_{\delta}$ (e.g., $v_{1}.UB_{3}$ and $v_{2}.UB_{2}$), and MBRs with same degree $q_{1}.MBR_{2}$ and $v_{i}.MBR_{2}$ (e.g., $v_{1}.MBR_{2}$ and $v_{2}.MBR_{2}$), respectively. After that, we can obtain the $q_{1}$’s candidate vertex set, $q_{1}.cand\_set$.
@@
Input: i) a query graph $q$; ii) a sorted list (query plan) $Q$; iii) a snapshot graph $G_{t}$; iv) matching results set AA; v) a sorted list, $M$, of vertices matching with query vertices in $Q$, and; vi) the recursion depth $n$
@@ 21  Refinement ($q,Q,G_{t},A,M,n+1$);
Specifically, we maintain a vertex vector $M$ that will store vertices of the subgraph $g$ matching with ordered query vertices in $Q$. To enumerate all matching subgraphs, each time we recursively expand partial matching results by including a new candidate vertex $u$ as $M[n]$ that maps with the $n$-th query vertex $Q[n]$. If we find all vertices in $M$ mapping with $Q$ (i.e., recursion depth $n=|Q|$), then we can add $M$ to the answer set $A$ (lines 1-3). Otherwise, we will find a vertex candidate set $S_{cand}$ from Q[n].cand_setQ[n].cand\_set, such that each vertex $u$ in $S_{cand}$ has the same edge connections to that in $M$ as that in $Q$ (lines 5-11). Next, for each vertex $u\in S_{cand}$, we treat it as vertex $M[n]$ matching with $Q[n]$, and recursively call function Refinement($\cdot$) with more depth $n+1$ (lines 13-17).
@@
For the synopsis traversal (lines 2-8), assume that $PP_{C}$ is the pruning power of the cells’ key value and $PP_{v}$ is the pruning power in each cell’s point list. Thus, the synopsis traversal cost is $O(|V(q)|\cdot(1-PP_{C})K^{d}\cdot(1-PP_{v})|C.list|)$, where $K^{d}$ is the number of cells and $|C.list|$ is the number of vertices in the cell $C$.
@@
Input: i) a dynamic graph $G_{t-1}$ at previous timestamp (t1)(t-1); ii) a graph update operation $\Delta G_{t}$; iii) a registered query graph $q$, and; iv) matching result set A(q,t1)A(q,t-1) at timestamp (t1)(t-1)
Output: an update matching result set, $A(q,t)$, at timestamp $t$
@@ 3  $\Delta A(q,t)\leftarrow\emptyset$
8           $M^{\prime}[0]=v_{i},M^{\prime}[1]=v_{j}$
9
[…]
@@ 14 $A(q,t)\leftarrow A(q,t-1)\cup\Delta A(q,t)$
  // edge deletion operation $\Delta G_{t}=(e-,t)$
17  A(q,t)=A(q,t1)A(q,t)=A(q,t-1)
@@ 20        $A(q,t)\leftarrow A(q,t)-\{g\}$
Algorithm 4 The Query Answering Algorithm for CSM Query
6.3. CSM Query Answering
Given a dynamic graph $G_{t-1}$ at timestamp (t1)(t-1), a registered query graph $q$, and a graph update operation $\Delta G_{t}$ at timestamp $t$, Algorithm 4 provides a CSM query answering algorithm, which maintains/obtains the latest subgraph matching answer set, $A(q,t)$. Specifically, if the graph update operation $\Delta G_{t}$ is an insertion of edge $e=(v_{i},v_{j})$, we will calculate an incremental update, $\Delta A(q,t)$, of new subgraph matching results (lines 1-11).
Edge Insertion: For the insertion of an edge $e=(v_{i},v_{j})$, we will first find matching edge(s) $(q_{i},q_{j})$ in the query graph with the same labels (line 3). Then, we will check whether or not edges $(v_{i},v_{j})$ and $(q_{i},q_{j})$ match each other, by applying the embedding dominance and MBR range pruning strategies (described in Section 6.1; lines 4-5). If the answer is yes, then we will generate a new query plan $Q^{\prime}$ starting from query vertices $q_{i}$ and $q_{j}$ (note: the remaining ones are obtained from $Q$, similar to line 10 of Algorithm 2; line 7), initialize the first two matching vertices $v_{i}$ and $v_{j}$ in a sorted list $M^{\prime}$ (lines 8-9), and invoke the function Refinement($\cdot$) with parameters of new query plan $Q^{\prime}$, incremental answer set $\Delta A(q,t)$, initialized sorted list $M^{\prime}$, and recursion depth $n=2$ (as two matching vertices have been found; line 10). After that, we can update the answer set $A(q,t)$ with $A(q,t-1)\cup\Delta A(q,t)$ (line 11).
Edge Deletion: When $\Delta G_{t}$ is an edge deletion operation (e,t)(e-,t), we can simply remove those existing subgraph matching answers $g\in A(q,t-1)$ which contain the deleted edge $e$, and obtain the updated answer set $A(q,t)$ (lines 12-16). Finally, we return the CSM query answer set $A(q,t)$ at current timestamp $t$ (line 17).
For the edge deletion operation (lines 12-16), since we delete the matching results from A(q,t1)A(q,t-1) that contain the deleted edge, by using the hash file to check the edge existence with $O(1)$ cost, the time complexity is given by O(|A(q,t1)|)O(|A(q,t-1)|).
Discussions on Batch Updates: Similar to existing works on CSM (Choudhury et al., 2015; Kankanamge et al., 2017; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021; Sun et al., 2022a), we adopt dynamic graph model (Definition 2.4) with one single edge update per timestamp. Thus, as a straightforward method, batch updates can be supported via sequential execution of atomic edge updates.
To enhance the efficiency of batch processing multiple graph edge updates at a time, we can group similar embeddings of query edges (from registered query graphs), as well as those of data edges affected by edge updates from the batch, into MBRs. Then, we can obtain candidate pairs of query/data MBRs, in which we filter/refine candidate query/data vertices. Specifically, to enable batch processing, in Algorithm 4, we can add an intersection checking between edge label MBRs for query/data edges (line 3). Then, we check the dominance between the minimum and maximum corner points of the query and data edge MBRs, respectively, and if the dominance holds, we consider candidate pairs of query/data vertices from MBR pairs (line 4). This way, we can save the cost of some unnecessary pairwise dominance checks for batch updates. We would like to leave the interesting topics of further optimizations for batch update as our future work.
Discussions on Dynamic Vertex Label Change: In this paper, our proposed DIVINE framework can naturally support dynamic vertex label changes by updating the embeddings of the vertex and its 1-hop neighbors. Specifically, for an affected vertex $v_{i}$, we only need to (1) update the SPUR vector of this vertex $v_{i}$, using seeded pseudo-random values derived from vertex labels, and; (2) update the SPAN vectors of 1-hop neighbors of vertex $v_{i}$, using $v_{i}$’s updated SPUR vector. Then, new embedding vectors of the affected vertices (i.e., $v_{i}$ and its 1-hop neighbors) will be updated in DAS3.
Discussions on Graphs with Directed, Labeled and/or Loop Edges: In this paper, our DIVINE approach focuses on a dynamic, undirected, vertex-label graph without loops. To consider labeled edges, we may construct the SPAN vector $y_{i}$ for each vertex $v_{i}$, by summing up SPUR vectors $x_{j}$ of $v_{i}$’s 1-hop neighbors $v_{j}$, as well as randomized vectors of edges $e_{i,j}$ (with their edge labels as seeds). To support CSM over graphs with directed and/or loop edges, we can first use DIVINE to compute candidate subgraphs by ignoring edge directions, and then apply a post-processing refinement phase that validates edge direction and/or loop constraints. We would like to leave the topics of generalizing our DIVINE method to graphs with directed, labeled, and/or loop edges as our future work.